OpenStudy (anonymous):

According to vander waal's equation prove that excluded volume = 4*actual volume i.e b=4v?

OpenStudy (anonymous):

Excluded volume of 2 molecules is sphere of radius of 2R where R radius of ...Excluded volume for two molecule Excluded volume for one molecule ? Constant 'a'measures force of attraction. Greater value 'a', higher intermolecular force of attraction.

OpenStudy (anonymous):

okay I got this : The equation can be cast into the better known form \left(p + \frac{n^2 a}{V^2}\right)\left(V-nb\right) = nRT where a = N_A^2a' is a measure of the attraction between the particles, b = N_A b' is the volume excluded by a mole of particles.

OpenStudy (anonymous):

i hav nt understood can u xplain it more

OpenStudy (anonymous):

Here this should help :D http://en.wikipedia.org/wiki/Van_der_Waals_equation